OpenStudy (andijo76):

I'm in high school and need help on Algebra Two homework. I need to know know how this is read: T:a-->a^2+3a+4. can anyone help??

OpenStudy (dumbcow):

this is a quadratic equation which graph is that of a parabola

OpenStudy (anonymous):

it is a fancy way to write function notation, often seen as \[f(x)=x^2+3x+4\]

OpenStudy (anonymous):

\(T\) sends \(a\) to \(a^2+3a+4\)

jimthompson5910 (jim_thompson5910):

T:a-->a^2+3a+4 means T is a function rule that maps 'a' to a^2+3a+4 essentially f(a) = a^2 + 3a + 4 is the function rule that you may be familiar with in terms of notation

OpenStudy (andijo76):

I need a little more detail like my first problem was f(x) and i had to write f of x. something like that. Would that be how you write it satellite73?

OpenStudy (anonymous):

i am not sure what the actual question is do you have something other than \[T:a\to a^2+3a+4\]?

OpenStudy (andijo76):

A(x)= 1/2x(5-x)

OpenStudy (anonymous):

i guess i am not clear my question is, what is your question ? what are you supposed to do with this?

OpenStudy (andijo76):

im supposed to tell how each is read

OpenStudy (anonymous):

"A of x equal one half x times the quantity five minus x "

OpenStudy (andijo76):

thank you.

OpenStudy (andijo76):

and the previous one?

OpenStudy (anonymous):

T send a to a squared plus 3 times a minus 4

OpenStudy (andijo76):

thank you

OpenStudy (anonymous):

yw

OpenStudy (triciaal):

* plus 4